Cooley Landscaping needs to borrow 35,000 for a new​ front-end dirt loader. The bank is willing to loan the money at 7.5​% interest for the next

7 years with annual,semiannual​,quarterly​,or monthly payments. What are the different payments that Cooley Landscaping could choose for these different payment​ plans?
Answer #1

1.Annual payment

Information provided:

Present value= $35,000

Time= 7 years

Interest rate= 7.5%

The amount of annual payment is calculated by entering the below in a financial calculator:

PV= -35,000

N= 7

I/Y= 7.5

Press the CPT key and PMT to compute the amount of annual payment.

The value obtained is 6,608.01.

Therefore, the amount of annual payment is $6,608.01.

2.Semi-annual payment

Information provided:

Present value= $35,000

Time= 7 years*2= 14 semi-annual periods

Interest rate= 7.5%/2= 3.75% per semi-annual period

The amount of semi-annual payment is calculated by entering the below in a financial calculator:

PV= -35,000

N= 14

I/Y= 3.75

Press the CPT key and PMT to compute the amount of semi-annual payment.

The value obtained is 3,258.96.

Therefore, the amount of semi-annual payment is $3,258.96.

3.Quarterly payment

Information provided:

Present value= $35,000

Time= 7 years*4= 28 quarters

Interest rate= 7.5%/4= 1.8750% per quarter

The amount of quarterly payment is calculated by entering the below in a financial calculator:

PV= -35,000

N= 28

I/Y= 1.8750

Press the CPT key and PMT to compute the amount of quarterly payment.

The value obtained is 1,618.13.

Therefore, the amount of quarterly payment is $1,618.13.

4.Monthly payment

Information provided:

Present value= $35,000

Time= 7 years*12= 84 months

Interest rate= 7.5%/12= 0.6250% per month

The amount of monthly payment is calculated by entering the below in a financial calculator:

PV= -35,000

N= 84

I/Y= 0.6250

Press the CPT key and PMT to compute the amount of monthly payment.

The value obtained is 536.84.

Therefore, the amount of monthly payment is $536.84.

In case of any query, kindly comment on the solution.